Project

General

Profile

Actions

Run Plan 2012-03-22

We will be running from Friday 23rd to Monday 26th of March with an option to extend to Tuesday 27th if there is a problem (e.g. ISIS down)

  • There will be no KL, TOF2, or decay solenoid during the run
  • We will have the following aims (and responsible person):
    • Test state machines and run control sequence - Hanlet
    • Test GRID/offline batch job - Rogers
    • Test cerenkov monitoring - Kafka/Cremaldi
    • Verify beam timing vs RF pulse - Overton
    • Test online monitoring visible through micewww - Rogers
    • D1 Scan - Rogers
    • (If time) - target tests - Overton/Smith

I plan initially about 4 hours for each activity. Given 2 hours to run up and down each shift, that means 4*8 hour shifts. Probably some things may be able to piggyback on other things and we can run for more than 8 hour shifts - but everything takes longer than we expect.

Shifters:

March... Th 22 Fr 23 Sa 24 Su 25 Mo 26 Tu 27
MOM Blackmore Blackmore Blackmore Blackmore Hodgson Hodgson
Lead shifter Hanlet Rogers Hanlet Hanlet Rogers Rogers
Shifter 1 Smith Smith Smith Smith Smith Smith
Shifter 2 Nebrensky Nebrensky Nebrensky Nebrensky Nebrensky Nebrensky
BLOC Overton Overton Overton Overton Overton Overton
SOC Rogers Rogers Rajaram Rajaram Rogers Rogers
Experimenter
Aim

SOC is software on call

Planning Meeting

We will have a planning meeting on 22nd March at 14.00-15.00 UK time in the Norton Room. Aim is to just go through the run plan and answer any questions.

Operations Issues

Jobs to do before the run starts

  • Paperwork (ATWs) <--- Done (Blackmore)
  • get cranes moved to W end of Hall (away from ISIS) <-- will be done Thursday morning (Blackmore)
  • Check PPS doors/BOBs OK -- Craig will be around Thursday afternoon to help with any PPS snags (Blackmore)
  • the phone seems to have disappeared from the desk by the DSA entrance <--- The phone is plugged in close to the DSA door, near the Linde fridge. As there are no free plug sockets on the desk by the DSA entrance, I assume that this is where it's meant to be. I think it used to live closer to the hall entrance. Given the lack of "Phone Here!" signs in the hall, they're always going to be difficult to find. (Blackmore)
  • Erect temporary fencing along Q789 and across gap by sliding floors <--- will be done Thursday morning (Blackmore)
  • Beamline polarity is currently -ve - probably want that changed for rate <----- Done on 21/03/12 (Blackmore)
  • Locate latest run summary spreadsheet - should have the DS inversion runs in it <--- Found 22/03, luckily it was on Henry's laptop...
  • turn SY527 HV supply on
  • make sure DAQ works (e.g. takes cosmic data) and is in a sensible configuration (selection of spill gate/trigger condition/event buffer etc.)
  • Check/fix MICERAID (see #945, #951 )

Updated by Nebrensky, Henry over 10 years ago ยท 39 revisions